Circular No. 6261

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

AX J2315–592 H.-C. Thomas, Max-Planck-Institut für Astrophysik, Garching; and K.

Reinsch, Universitäts-Sternwarte Göttingen, write: “We have identified the newly-discovered x-ray source AX J2315–592 (IAUC 6260) as a new AM Her binary. The star (V = 17.1) located 15′′ north of the x-ray position at α = 23^{h}15^{m}18^{s}.4, δ = −59^{o}10′27′′ (equinox 2000.0) has been observed with the ESO/MPI 2.2-m telescope (+ EFOSC2) at La Silla on Nov. 12 and 13. The spectra display strong emission lines of the Balmer series and of helium. A radial-velocity variation with an amplitude of 200 km/s and an intensity variation between mag 14.4 and 16.2 in the Gunn i filter have been found with a period of 5360 ± 10 s. The variation in brightness decreases toward shorter wavelengths, indicating strong cyclotron radiation to be present at long wavelengths. No cyclotron harmonics could be distinguished, and this must be due to a rather low magnetic field strength at the white dwarf’s surface. At certain phases, the emission lines have extended blue wings. No features of the M star have been detected in the spectra.”

51 PEGASI E. Guinan, Villanova University, writes: “High-precision photoelectric

photometry of 51 Peg has been carried out at Mount Hopkins by R. Dukes, H. Nations, and D. Buzasi (College of Charleston) and by E. Guinan and G. McCook since Oct. 14. No significant light variations have been detected for 51 Peg to the level of the photometric precision (0.0018 mag). Some of the photometry was carried out near the expected times of the companion’s transit eclipse. Several nearby comparison and check stars have been used for this differential photometry. As a cautionary note to other observers, one of the most logical primary comparison stars for the photometric study of 51 Peg, HD 217813 (V = 6.60, B − V = +0.61; spectral type G5 V), was found by both groups to be a variable star, showing quasi-sinusoidal light variations with a period of ∼ 8.5 days and brightness ranges of 0.012 mag in V and 0.020 mag in U; the period and wavelength dependence are consistent with the rotational modulation in brightness by starspots from a Hyades-age solar-type star. Another possible photometric reference star near the position of 51 Peg, HD 218396 (= HR 8799; A7–F0 V; V = 5.99, B − V = +0.25), is also a small-amplitude variable (see Rodriguez and Zerbi, IBVS No. 4170). HD 217924 (V = 7.3, B − V = +0.6; G0 V) shows no indications of photometric variations greater than ∼ 0.002 mag and seems to be suitable as a comparison star for the study of 51 Peg.”

1995 November 14 Daniel W. E. Green